OpenStudy (anonymous):

The ratio of concentrate to water is listed for each punch below. Punch A: 1:6 Punch B: 2:9 Punch C: 1:3 Punch D: 5:12 Which ratio gives the strongest punch?

OpenStudy (domebotnos):

so its water to punch right?

OpenStudy (anonymous):

yes, I think so idk im not smart

OpenStudy (here_to_help15):

Turn this into a fraction @emilybunny42 actually turn all of them into fractions

OpenStudy (domebotnos):

nooo you are smart! what do you think it is??

OpenStudy (anonymous):

let me think...

OpenStudy (anonymous):

uhhhhhhhhhh idk, see I hate math, and all of da numbers

OpenStudy (welshfella):

convert the rations to concentration of punch as a decimal eg for A it is 1 / (1+6) = 1/7 = 0.1429 do this for every Punch and the answer will be obvious

OpenStudy (domebotnos):

MY opinion is A, now im not saying that is the answer, but that just seems more logical to me...

OpenStudy (here_to_help15):